# Neyman-Pearson multiclass classification under label noise via empirical likelihood

### 23 Mar 2026 v1

@@ 4.1 Revisiting NP binary classification with clean labels
$\begin{split}\min_{\phi}&~P_{1}^{*}(\{\phi(X)\neq 1\})\\ \text{s.t.}&~P_{0}^{*}(\{\phi(X)\neq 0\})\leq\alpha,\end{split}$ (22)
where $\alpha\in(0,1)$ is the desired type I error level, and the objective represents the type II error. This formulation is a special case of the general NPMC problem in (1) with $\rho_{0}=0$, $\rho_{1}=1$, $\mathcal{S}=\{0\}$, and $\alpha_{0}=\alpha$.
When the joint distribution $P_{X,Y}^{*}$ is known, the NP lemma (Sadinle et al., 2019, Lemma 1) establishes that the optimal classifier takes the form:
@@ 4.2 NP binary classification with noisy labels with statistical guarantees
Let $\widehat{\phi}_{\widehat{\lambda}}(x)$ be the binary classifier defined in (27). Under Assumptions 4.1 and 4.2, the type I error is controlled as follows:
$P_{0}^{*}\left(\left\{\widehat{\phi}_{\widehat{\lambda}}(X)\neq 0\right\}\right)\leq\alpha+O_{p}(n^{-1/2}).$
@@
The NP umbrella algorithm was developed by Tong et al. (2018) to control errors in binary classification with uncontaminated data, and later extended to the noisy-label setting by Yao et al. (2023). When domain knowledge of the noise transition matrix $\bm{M}$$^{*}$ is available, they estimate the discrepancy between true and corrupted type I errors and propose a label-noise-adjusted NP umbrella algorithm with theoretical guarantees. Specifically, for any $\delta>0$, they show that their classifier $\phi_{\delta}$ satisfies
@@
where $\varepsilon(n)$ vanishes as $n$ increases. To our knowledge, this is the only work addressing error control under noisy labels. It is therefore natural to compare our error control guarantee (Theorem 4.1) with their result. At first glance, our guarantee appears weaker than that of Yao et al. (2023). In particular, we show that $\forall\epsilon>0$, there exist constants $C$ and $N$ such that for all $n\geq N$,
@@
[…]
@@ 4.3 Simulation study
In this section, we evaluate the performance of our proposed classifier (27) on simulated data and compare it with existing methods. Following the experimental settings of Yao et al. (2023), we consider three distinct distributions for the feature variables $X^{y}$:
@@
For our method, we use linear basis functions in all experiments. These three cases are designed to evaluate both efficiency and robustness: the DRM assumption holds in Case A, whereas the model is misspecified in Cases B and C. We compare our method with the Vanilla and NPC algorithms proposed in Yao et al. (2023).
@@ 5 Neyman-Pearson multiclass classification
In this section, we study the more general NPMC problem. Extending NP binary classification to the multiclass setting is non-trivial for two main reasons as given in Tian and Feng (2025): (i) the NPMC problem may not always be feasible, and (ii) solving it requires a more involved dual-form numerical procedure. To address this, we adopt their NPMC solver and combine it with our EL-based consistent estimators of $\{w_{k}^{*},\pi_{k}^{*}(x)\}_{k=0}^{K-1}$ to handle NPMC under noisy labels. When feasibility holds, we establish that the resulting classifiers control class-specific type I errors up to a term that vanishes as the sample size increases. Empirically, we show that our method performs on par with the clean-label benchmark when the sample size is sufficiently large.
In this section, we first review their NPMC solver under clean data, then introduce our estimator along with its theoretical guarantees, and finally present the empirical results.
@@ 5.1 Revisiting NPMC with clean labels
Tian and Feng (2025) solves (1) via its dual formulation. The Lagrangian associated with (1) is
@@
where $\mbox{$\lambda$}=\{\lambda_{k}\}_{k\in{\mathcal{S}}}\in{\mathbb{R}}_{+}^{|{\mathcal{S}}|}$ are the Lagrange multipliers. For any fixed $\lambda$, the optimal classifier that minimizes the Lagrangian has the closed form (Tian and Feng, 2025, Lemma 2):
@@
with $c_{k}(\mbox{$\lambda$},{\bm{w}})=\{\rho_{k}+\lambda_{k}\mathbbm{1}(k\in{\mathcal{S}})\}/w_{k}$. If strong duality holds, the optimal $\lambda$ can be obtained by solving the dual problem:
@@
$G(\mbox{$\lambda$})=$ $\min_{\phi}L(\phi,\mbox{$\lambda$})=L(\phi^{*}_{\mbox{$\lambda$}},\mbox{$\lambda$})$
$=$ $-{\mathbb{E}}_{X}\left\{c_{\phi^{*}_{\mbox{$\lambda$}}(X)}(\mbox{$\lambda$},{\bm{w}}^{*}){\mathbb{P}}(Y=\phi^{*}_{\mbox{$\lambda$}}(X)|X)\right\}+\sum_{k\in[K]}\rho_{k}+\sum_{k\in{\mathcal{S}}}\lambda_{k}(1-\alpha_{k}).$ (32)
The key advantage of the dual formulation is that the dual objective $G(\mbox{$\lambda$})$ is concave, regardless of whether the primal problem is convex. This concavity allows us to first solve for $\mbox{$\lambda$}^{*}$ using a standard convex optimization method, and then substitute $\mbox{$\lambda$}^{*}$ into (30) to obtain the optimal classifier $\phi^{*}_{\mbox{$\lambda$}^{*}}$.
Given a finite training set ${\mathcal{D}}^{\text{train}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$Tian and Feng (2025) proposes the NPMC-CX (ConveX) estimator for the dual objective $G(\mbox{$\lambda$})$:
@@
For readers less familiar with optimization, we briefly explain the motivation for solving the dual problem and its connection to the primal problem. The primal problem (1) involves constraints on class-specific error for certain classes, which can make direct optimization challenging. By introducing Lagrange multipliers $\lambda$, we form the Lagrangian (29), and define the dual function $G(\mbox{$\lambda$})$ as the minimum of the Lagrangian over all classifiers. These two optimization problems are connected via the weak duality property, which states that for any feasible $\mbox{$\lambda$}\geq 0$, the dual objective $G(\mbox{$\lambda$})$ provides a lower bound on the primal objective. This ensures that maximizing the dual gives a value that never exceeds the optimal value of the primal. If strong duality holds, meaning that the maximum of the dual objective equals the minimum of the primal, solving the dual problem yields the optimal multipliers $\mbox{$\lambda$}^{*}$, which can be substituted into (30) to recover the primal optimal classifier. The key advantage of working with the dual is that $G(\mbox{$\lambda$})$ is concave regardless of the convexity of the primal, making the problem easier to solve with standard convex optimization techniques. We refer interested readers to Boyd and Vandenberghe (2004, Chapter 5) for more details on duality.
@@ 5.2 NPMC with noisy labels with statistical guarantees
Specifically, let $\widehat{{\bm{w}}}$, $\widehat{\mbox{$\gamma$}}$, and $\widehat{\mbox{$\beta$}}$ denote the MELE estimators in (14) for $\bm{w}$$^{*}$, $\mbox{$\gamma$}^{*}$, and $\mbox{$\beta$}^{*}$, respectively. Define
@@
To construct the final EL-CX classifier, in addition to computing $\widehat{\bm{w}}$ and $\widehat{\pi}_{k}(x)$ via the EM algorithm, one also needs to maximize $\widehat{G}(\mbox{$\lambda$})$ to obtain $\widehat{\mbox{$\lambda$}}$. We perform this maximization using the Hooke–Jeeves algorithm proposed in Tian and Feng (2025) (implemented in pymoo Blank and Deb 2020). We adopt the default initialization strategy in pymoo to generate the initial search points. The optimization is constrained to the hypercube $0,200]^{|{\mathcal{S}}|}$.
@@
Let $\mbox{$\lambda$}^{*}$ and $G(\mbox{$\lambda$})$ as defined in (31) and (32) respectively. We assume that $G(\mbox{$\lambda$})$ is twice continuously differentiable and $\nabla^{2}G(\mbox{$\lambda$}^{*})\prec 0$.
@@
For any classifier $\phi$, denote the $k$th class-specific misclassification probability as $R_{k}(\phi)={\mathbb{P}}(\phi(X)\neq k|Y=k)$. Let $\widehat{\phi}_{\widehat{\mbox{$\lambda$}}}$ be our EL-CX based classifier in (35). When the NPMC problem (1) is feasible, under Assumptions 4.25.1, and 5.2, we have that for any $k\in{\mathcal{S}}$:
$R_{k}(\widehat{\phi}_{\widehat{\mbox{$\lambda$}}})\leq\alpha_{k}+O_{p}(\varepsilon(n)),$
where $\varepsilon(n)\to 0$ as $n\to\infty$.
[…]
@@
Tian and Feng (2025) also proposes an alternative estimator for $G(\mbox{$\lambda$})$ based on a sample-splitting strategy. While this estimator relaxes the assumption on the Rademacher complexity of the function class $\{\pi_{k}(x)\}$, it has lower efficiency due to sample splitting. As shown in the proof of Theorem 5.1, our model satisfies the Rademacher complexity condition, so we do not adopt this alternative estimator in our work.
@@ 5.3 Simulation study
In this section, we conduct empirical experiments to compare our proposed method with other approaches. Following Tian and Feng (2025), we consider three typical settings to generate data:
@@
Under each case, we generate a training set of size $n$ and a test set of fixed size m=20,000m=20{,}000. Since we study the NPMC problem under noisy labeled data, we generate noisy labels for the training set using the following transition matrix:
$T_{lk}^{*}=\begin{cases}\eta/K&\text{if }k\neq l,\\ 1-(K-1)\eta/K&\text{if }k=l.\end{cases}$
We vary $n$ from $5000$ to $25,000$ with a step size of $5,000$, and consider $\eta\in\{0.05,0.1,0.15,0.2\}$. Each setting is repeated $R=500$ times.
@@
We compare EL-CX with the following baselines: (1) Oracle, which applies NPMC-CX (Tian and Feng, 2025) to clean training labels; and (2) Vanilla, which applies NPMC-CX directly to noisy labels without accounting for label noise.
[…]
@@
[…]
@@ 6 Real data experiments
[…]
@@ 7 Discussion and concluding remarks
Incorporating data-adaptive selection of the basis function $g$ as in Zhang and Chen (2022), or even integrating it with deep neural networks, remains an important direction for future work. In addition, extending our framework to handle more complex instance-dependent label noise (Xia et al., 2020) is another promising yet challenging avenue.
Acknowledgement
The authors would like to thank Yan Shuo Tan for insightful discussions. Qiong Zhang gratefully acknowledges the funding support by the National Key R &D Program of China Grant 2024YFA1015800.
@@ Acknowledgement
[…]
@@
    Note that $Q$ is separable in $\bm{w}$}, $\bm{T}$}, and $(\mbox{$\gamma$},\mbox{$\beta$})$. Thus, we can optimize each set of parameters separately, as described below. We now present the details of the M-step with respect to the three sets of variables.
@@ 11 $(1,g^{\top}(X_{n}))^{\top}$ $\omega_{n1}^{(t)}$
$T_{kk}^{*}\geq\xi_{k},\quad\xi_{k}\in[0,1],\quad k=1,\ldots,K$
@@
[…]
@@
[…]
@@
[…]
In this section, we discuss how to integrate our EL-based estimator under noisy labels with the framework of Tong et al. (2018) to achieve error control in binary classification.
The NP umbrella algorithm, proposed by Tong et al. (2018), provides a flexible procedure that accommodates various score functions from binary classifiers. It produces a classifier that controls the type I error at the target level with high probability while minimizing the type II error. The procedure is as follows:

### 21 Apr 2026 v2

@@ 4.1 Revisiting NP binary classification with clean labels
$\min_{\phi}~~P_{1}^{*}(\{\phi(X)\neq 1\})\qquad\qquad\text{s.t.}~~P_{0}^{*}(\{\phi(X)\neq 0\})\leq\alpha,$ (21)
where $\alpha\in(0,1)$ is the target type I error, and the objective corresponds to the type II error. This is a special case of (1) with $\rho_{0}=0$, $\rho_{1}=1$, ${\mathcal{S}}=\{0\}$, and $\alpha_{0}=\alpha$.
When the joint distribution $P_{X,Y}^{*}$ is known, the NP lemma (Sadinle et al., , 2019, Lemma 1) establishes that the optimal classifier takes the form:
@@ 4.2 NP binary classification with noisy labels with statistical guarantees
Let $\widehat{\phi}_{\widehat{\lambda}}(x)$ be the binary classifier defined in (26). Under Assumptions 4.1 and 4.2, the type I error is controlled as: $P_{0}^{*}\left(\left\{\widehat{\phi}_{\widehat{\lambda}}(X)\neq 0\right\}\right)\leq\alpha+O_{p}(n^{-1/2})$.
@@
The NP umbrella algorithm was developed by Tong et al., (2018) to control errors in binary classification with uncontaminated data, and later extended to the noisy-label setting by Yao et al., (2023). When domain knowledge of the noise transition matrix $\bm{M}$$^{*}$ is available, they estimate the discrepancy between true and corrupted type I errors and propose a label-noise-adjusted NP umbrella algorithm with theoretical guarantees. Specifically, for any $\delta>0$, they show that their classifier $\phi_{\delta}$ satisfies
@@
where $\varepsilon(n)$ vanishes as $n$ increases. To our knowledge, this is the only work addressing error control under noisy labels. It is therefore natural to compare our error control guarantee (Theorem 4.1) with their result. At first glance, our guarantee appears weaker than that of Yao et al., (2023). In particular, we show that $\forall\epsilon>0$, there exist constants $C$ and $N$ such that for all $n\geq N$,
@@
Comparing this with (27), the distinction lies in the quantities being controlled. Our analysis bounds the deviation of the empirical type I error from the target level, which introduces a stochastic fluctuation of order $n^{-1/2}$. In contrast, the NP umbrella framework provides a stricter guarantee, ensuring with high probability that the population type I error never exceeds the target level. Importantly, our EL-based estimator is flexible and can be paired with different error control methods: when combined with the NP umbrella algorithm, it yields the same guarantee as in Yao et al., (2023). Moreover, our method does not require prior knowledge of $\bm{M}$$^{*}$. Further discussion and empirical results are provided in Appendix B.2.3.
@@ 4.3 Simulation study
In this section, we evaluate the performance of our proposed classifier (26) on simulated data and compare it with existing methods. Following the experimental settings of Yao et al., (2023), we consider three distinct distributions for the feature variables $X^{y}$:
@@
For our method, we use linear basis functions in all experiments. These three cases are designed to evaluate both efficiency and robustness: the DRM assumption holds in Case A, whereas the model is misspecified in Cases B and C. We compare our method with the Vanilla and NPC algorithms proposed in Yao et al., (2023).
@@ 5 Neyman-Pearson multiclass classification
In this section, we study the more general NPMC problem. Extending NP binary classification to the multiclass setting is nontrivial for two reasons (Tian and Feng, , 2025): (i) the NPMC problem may be infeasible, and (ii) solving it requires a more involved dual formulation. We adopt their NPMC solver and combine it with our EL-based estimators of $\{w_{k}^{*},\pi_{k}^{*}(x)\}_{k=0}^{K-1}$ to handle noisy labels. When feasible, we show that the resulting classifiers control class-specific type I errors up to a vanishing term. Empirically, our method matches the clean-label benchmark with sufficiently large samples. For completeness, we first review their solver under clean data, then introduce our estimator with theoretical guarantees, and finally present empirical results.
@@ 5.1 Revisiting NPMC with clean labels
Tian and Feng, (2025) solves (1) via its dual formulation. The Lagrangian associated with (1) is
@@
where $\mbox{$\lambda$}=\{\lambda_{k}\}_{k\in{\mathcal{S}}}\in{\mathbb{R}}_{+}^{|{\mathcal{S}}|}$ are the Lagrange multipliers. For any fixed $\lambda$, the optimal classifier that minimizes the Lagrangian has the closed form (Tian and Feng, , 2025, Lemma 2):
@@
with $c_{k}(\mbox{$\lambda$},{\bm{w}})=\{\rho_{k}+\lambda_{k}\mathbbm{1}(k\in{\mathcal{S}})\}/w_{k}$.
If strong duality holds, the optimal $\lambda$ can be obtained by solving the dual problem:
@@
$\begin{split}G(\mbox{$\lambda$})=&\min_{\phi}L(\phi,\mbox{$\lambda$})=L(\phi^{*}_{\mbox{$\lambda$}},\mbox{$\lambda$})\\ =&-{\mathbb{E}}_{X}\left\{c_{\phi^{*}_{\mbox{$\lambda$}}(X)}(\mbox{$\lambda$},{\bm{w}}^{*}){\mathbb{P}}(Y=\phi^{*}_{\mbox{$\lambda$}}(X)|X)\right\}+\sum_{k\in[K]}\rho_{k}+\sum_{k\in{\mathcal{S}}}\lambda_{k}(1-\alpha_{k}).\end{split}$
The key advantage of the dual formulation is that the dual objective $G(\mbox{$\lambda$})$ is concave, regardless of whether the primal problem is convex. This allows us to solve for $\mbox{$\lambda$}^{*}$ via standard convex optimization, and then substitute it into (29) to obtain the optimal classifier $\phi^{*}_{\mbox{$\lambda$}^{*}}$.
Given a finite training set ${\mathcal{D}}^{\text{train}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$Tian and Feng, (2025) proposes the NPMC-CX (ConveX) estimator for the dual objective $G(\mbox{$\lambda$})$:
@@
For readers less familiar with optimization, we briefly explain the motivation for solving the dual problem and its connection to the primal problem. The primal problem (1) involves constraints on class-specific error for certain classes, which can make direct optimization challenging. By introducing Lagrange multipliers $\lambda$, we form the Lagrangian (28), and define the dual function $G(\mbox{$\lambda$})$ as the minimum of the Lagrangian over all classifiers. These two optimization problems are connected via the weak duality property, which states that for any feasible $\mbox{$\lambda$}\geq 0$, the dual objective $G(\mbox{$\lambda$})$ provides a lower bound on the primal objective. This ensures that maximizing the dual gives a value that never exceeds the optimal value of the primal. If strong duality holds, meaning that the maximum of the dual objective equals the minimum of the primal, solving the dual problem yields the optimal multipliers $\mbox{$\lambda$}^{*}$, which can be substituted into (29) to recover the primal optimal classifier. The key advantage of working with the dual is that $G(\mbox{$\lambda$})$ is concave regardless of the convexity of the primal, making the problem easier to solve with standard convex optimization techniques. We refer interested readers to Boyd and Vandenberghe, (2004, Chapter 5) for more details on duality.
@@ 5.2 NPMC with noisy labels with statistical guarantees
Let $\widehat{{\bm{w}}}$, $\widehat{\mbox{$\gamma$}}$, and $\widehat{\mbox{$\beta$}}$ denote the MELE estimators in (13) for $\bm{w}$$^{*}$, $\mbox{$\gamma$}^{*}$, and $\mbox{$\beta$}^{*}$, respectively. Define
@@
To construct the final EL-CX classifier, in addition to computing $\widehat{\bm{w}}$ and $\widehat{\pi}_{k}(x)$ via the EM algorithm, one also needs to maximize $\widehat{G}(\mbox{$\lambda$})$ to obtain $\widehat{\mbox{$\lambda$}}$. We perform this maximization using the Hooke–Jeeves algorithm proposed in Tian and Feng, (2025) (implemented in pymoo Blank and Deb, 2020). We adopt the default initialization strategy in pymoo to generate the initial search points. The optimization is constrained to the hypercube $0,200]^{|{\mathcal{S}}|}$.
@@
Let $\mbox{$\lambda$}^{*}$ and $G(\mbox{$\lambda$})$ as defined in (30) and (5.1) respectively. We assume that $G(\mbox{$\lambda$})$ is twice continuously differentiable and $\nabla^{2}G(\mbox{$\lambda$}^{*})\prec 0$.
@@
For any classifier $\phi$, denote the $k$th class-specific misclassification probability as $R_{k}(\phi)={\mathbb{P}}(\phi(X)\neq k|Y=k)$. Let $\widehat{\phi}_{\widehat{\mbox{$\lambda$}}}$ be our EL-CX based classifier in (33). When the NPMC problem (1) is feasible, under Assumptions 4.25.1, and 5.2, we have that for any $k\in{\mathcal{S}}$: $R_{k}(\widehat{\phi}_{\widehat{\mbox{$\lambda$}}})\leq\alpha_{k}+O_{p}(\varepsilon(n))$, where $\varepsilon(n)\to 0$ as $n\to\infty$.
The proof is in Appendix C.7. This theorem establishes that when the primal problem is feasible, our classifier asymptotically controls the class-specific errors of the target classes at the target level. Both Assumptions 5.1 and 5.2 are the same as that in Tian and Feng, (2025).
@@
Tian and Feng, (2025) also proposes an alternative estimator for $G(\mbox{$\lambda$})$ based on a sample-splitting strategy. While this estimator relaxes the assumption on the Rademacher complexity of the function class $\{\pi_{k}(x)\}$, it has lower efficiency due to sample splitting. As shown in the proof of Theorem 5.1, our model satisfies the Rademacher complexity condition, so we do not adopt this alternative estimator in our work.
@@ 5.3 Simulation study
In this section, we conduct empirical experiments to compare our proposed method with other approaches. Following Tian and Feng, (2025), we consider three typical settings to generate data:
@@
Under each case, we generate a training set of size $n$ and a test set of fixed size m=20,000m=20{,}000. Since we study the NPMC problem under noisy labeled data, we generate noisy labels for the training set using the following transition matrix: $T_{lk}^{*}=\eta/K$ if $k\neq l$ and $1-(K-1)\eta/K$ otherwise. We vary $n$ from $5000$ to $25,000$ with a step size of $5,000$, and consider $\eta\in\{0.05,0.1,0.15,0.2\}$. Each setting is repeated $R=500$ times.
@@
We compare EL-CX with the following baselines: (1) Oracle, which applies NPMC-CX (Tian and Feng, , 2025) to clean training labels; and (2) Vanilla, which applies NPMC-CX directly to noisy labels without accounting for label noise.
As shown in Appendix B.1, the DRM assumption (6) holds in all cases with $g(x)=x$ for the first two cases and $g(x)=(x^{\top},x^{\top}x)^{\top}$ for the third case. We use the same covariates in the multinomial logistic regression model for all methods. The MSE of our MELE estimator converges at the root-$n$ rate, visualized in Figure 2 in Appendix B.2.1.
@@
We also provide violin plots of the class-specific error and the objective value under different noise levels $\eta$ for case (a) in Figure 1. The corresponding plots for the other two cases are included in Appendix B.2.2.
@@ 6 Real data experiments
Following Tian and Feng, (2025), we conduct experiments on two real datasets: the Dry Bean dataset and the Landsat Satellite dataset. Both datasets are used for multi-class classification. Due to space constraints, we present the results for the Landsat Satellite dataset here, and defer the results for the Dry Bean dataset to Appendix B.2.4.
@@ 7 Discussion and concluding remarks
Incorporating data-adaptive selection of the basis function $g$ as in Zhang and Chen, (2022), or even integrating it with deep neural networks, remains an important direction for future work. In addition, extending our framework to handle more complex instance-dependent label noise (Xia et al., , 2020) is another promising yet challenging avenue.
@@ A.1 EM algorithm implementation
where the maximization is subject to the constraints in (10). Using the same derivation from Appendix C.2, the profile complete data log-EL becomes
@@
    Note that $Q$ is separable in $\bm{w}$}, $\bm{T}$}, and $(\mbox{$\gamma$},\mbox{$\beta$})$. Since it can be written as an additive sum:
@@ A.3 Incorporating prior knowledge on transition matrix
$T_{kk}^{*}\geq\xi_{k},\qquad\xi_{k}\in[0,1],\qquad k=1,\ldots,K$
@@ A.4 Penalized EM algorithm
Appendix B Experimental details and more results
B.1 Density ratio model under normal model
@@
B.2 More experiment results
B.2.1 Model fitting under noisy labeled data
@@ B.2.2 Violin plots for case (b) and (c)
B.2.3 NP umbrella type control
In this section, we discuss how to integrate our EL-based estimator under noisy labels with the framework of Tong et al., (2018) to achieve error control in binary classification.
The NP umbrella algorithm, proposed by Tong et al., (2018), provides a flexible procedure that accommodates various score functions from binary classifiers. It produces a classifier that controls the type I error at the target level with high probability while minimizing the type II error. The procedure is as follows: